A mass 10kg is dropped from a height of 50cm.Find its
a)Potential Energy just before dropping.
b)Kinetic Energy just on touching the ground.
c)Velocity with which it hits the ground (given g=10ms-2)
1
Expert's answer
2012-02-24T08:27:41-0500
a) Potential energy
Ep=mgh=10kg*10m/s2*0.5m=50 J
b) Due to conservation of energy law
Ek=Ep=50 J
c) We know that Ek=mv2/2 hence we can find velocity
v=sqrt(2*Ek/m)=sqrt(2*50/10)=sqrt(10)=3,16 m/s